(The Birthday Problem) It is known that there are as many as n people in a room
a. How big is the chance that no one has the same birthday?
b. What is the value of n so that the probability of problem (a) is less than 50%?
Answer #1

a)P( no one has the same birthday )=P(all n people have different birthdays)=P(selecting n days from 365)

=(365*364*363*.....(365-n+1)/365n

b) P( no one has the same birthday )=(365*364*363*.....(365-n+1)/365n <0.50

solving above by putting values , n=23